OpenStudy (anonymous):

Im sorry LCM of 275x and 325x3

OpenStudy (anonymous):

factor them both into their primes. The lcm is the lowest power of each prime in both ex for lcm(25,125) 25 = 5^2 125=5^3 lowest power is 2 therefore lcm = 5^2

OpenStudy (anonymous):

also for 25 and 35 25 = 5^2 7^0 35 = 5^1 7^1 therefore lcm = 5^1 and 7^0 or just 5